determine the six circular function pi Over 4

Answer by jim_thompson5910(35256)

Using the unit circle, we get

sin(pi/4) = sqrt(2)/2

cos(pi/4) = sqrt(2)/2

tan(pi/4) = 1

csc(pi/4) = sqrt(2)

sec(pi/4) = sqrt(2)

cot(pi/4) = 1